# Angular integrals with three denominators via IBP, mass reduction, dimensional shift, and differential equations

### v1

@@ VIII.2 Single-massive integral
The quantities $Y_{123}^{(0,1)}$ are given in eqs. (39) respectively (40). As in the massless case, the result simplifies to a sum of two-denominator integrals in the case where the vectors $\vec{v}_{1,2,3}$ become linearly dependent. The pole part is identical to the result found in [7], while the finite part agrees in the limit of small mass $v_{11}$ with the result calculated in [7] via expansion by regions.
VIII.3 Double- and triple-massive integral
The $\varepsilon$-expansions of the double- and triple-massive integral could be obtained by plugging the massless and single-massive results from eqs.(66) and (73) into eqs.(14) and (15). However, it is much more economical to first use eq. (24) to extract the part that is expressible in terms of known two-denominator integrals. Subsequently eqs.(14) and (15) need to be applied only to $I_{111}^{d=6}$. By absorbing a factor of $\sqrt{Y_{123}}$, we can apply the mass splitting in the form of (21) respectively (22). Therefore the order $\varepsilon$ term in the double- respectively triple-massive case is conveniently expressed in terms of the functions
$f_{Y}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=f_{Y}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)-f_{Y}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)-f_{Y}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right),$ (78)
@@ VIII.3 Double- and triple-massive integral
$f_{Y}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=f_{Y}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)-f_{Y}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)$
$+f_{Y}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)+f_{Y}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)-f_{Y}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)$
@@
$I_{111}^{(2)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{1}{v_{13}v_{23}}+\frac{\pi}{2v_{12}v_{23}v_{13}-v_{11}v_{23}^{2}-v_{22}v_{13}^{2}}$
×$\times\left[\Big{(}v_{13}-v_{12}-v_{23}+\frac{v_{11}v_{23}}{v_{13}}\Big{)}\log\frac{v_{11}}{v_{13}^{2}}\right.+\Big{(}v_{23}-v_{12}-v_{13}+\frac{v_{22}v_{13}}{v_{23}}\Big{)}\log\frac{v_{22}}{v_{23}^{2}}$
@@
$+\frac{\pi}{2v_{12}v_{23}v_{13}-v_{11}v_{23}^{2}-v_{22}v_{13}^{2}}\left[\frac{1}{\sqrt{X_{12}}}\Big{(}v_{12}\left(v_{13}+v_{23}-v_{12}\right)-v_{11}v_{23}-v_{22}v_{13}+v_{11}v_{22}\Big{)}f_{1}^{(2)}(v_{12},v_{11},v_{22})\right.$
@@
$\left.\qquad+2\sqrt{Y_{123}^{(2)}}\,f_{Y}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})\right]+\mathcal{O}\!\left(\varepsilon^{2}\right)\,.$ (83)
with $f_{Y}^{(2)}$ from eq. (78). Finally, using the kinematic quantities from section IV, the triple-massive three-denominator integral has the $\varepsilon$-expansion
$I_{111}^{(3)}$ $=\frac{\pi}{X_{123}^{(3)}}\left[\frac{X_{12\bar{3}}}{\sqrt{X_{12}}}\log\!\left(\frac{v_{12}+\sqrt{X_{12}}}{v_{12}-\sqrt{X_{12}}}\right)+\frac{X_{1\bar{2}3}}{\sqrt{X_{13}}}\log\!\left(\frac{v_{13}+\sqrt{X_{13}}}{v_{13}-\sqrt{X_{13}}}\right)+\frac{X_{\bar{1}23}}{\sqrt{X_{23}}}\log\!\left(\frac{v_{23}+\sqrt{X_{23}}}{v_{23}-\sqrt{X_{23}}}\right)\right]$
$+\frac{\pi\,\varepsilon}{X_{123}^{(3)}}\left[\frac{X_{12\bar{3}}}{\sqrt{X_{12}}}f_{1}^{(2)}(v_{12},v_{11},v_{22})+\frac{X_{1\bar{2}3}}{\sqrt{X_{13}}}f_{1}^{(2)}(v_{13},v_{11},v_{33})+\frac{X_{\bar{1}23}}{\sqrt{X_{23}}}f_{1}^{(2)}(v_{23},v_{22},v_{33})+2\sqrt{Y_{123}^{(3)}}\,f_{Y}^{(3)}\right]+\mathcal{O}\!\left(\varepsilon^{2}\right)$ (84)
with $f_{Y}^{(3)}$ from eq. (79). For the double- and triple-massive integral, poles and small-mass limit of the finite part are also in agreement with the results from [7].
IX Beyond three denominators
If we consider a situation in $d=4-2\varepsilon$ dimensions where all vectors $v_{i}$ are confined to the physical $4$-dimensional subspace – an assumption applicable if the $v_{i}$ correspond to momenta of observed particles – we can reduce angular integrals with an arbitrary number of denominators raised to integer powers to the three-denominator case presented in this paper. This is possible since in $4$ dimensions, i.e. three spatial dimensions, no more than three $\vec{v}_{i}$ can be linearly dependent. Hence for any $n>3$ and vectors $v_{1},\dots,v_{n}$ there are constants $\lambda_{i}$ such that $\lambda\equiv\sum_{i=1}^{n}\lambda_{i}\neq 0$ but $0=\sum_{i=1}^{n}\lambda_{i}\vec{v}_{i}$. By partial fractioning it holds that [5]
@@ X Conclusion
In parallel work, Taushif Ahmed, Syed Mehedi Hasan, and Andreas Rapakoulias independently calculated the angular integral with three denominators using Mellin-Barnes integrals [52]. Preliminary numerical comparison at selected phase space points showed agreement.
@@ Appendix A Scalar products for mass reduction formula
The scalar products involving the massless auxiliary vectors appearing in section III are given by ($i,j,k\in\{1,2,3\}$)
@@
$v_{(ij)(ik)}$ $=\frac{1}{(2v_{ij}-v_{ii}-v_{jj})(2v_{ik}-v_{ii}-v_{kk})}\Big{[}v_{ii}^{2}v_{jk}-v_{ii}v_{ij}v_{ik}-v_{ii}v_{ij}v_{jk}-v_{ii}v_{ik}v_{jk}+v_{ii}v_{jj}v_{kk}+v_{ij}v_{ik}^{2}$
$\qquad+v_{ij}^{2}v_{ik}-v_{jj}v_{ik}^{2}+v_{ij}v_{ik}v_{jk}-v_{kk}v_{ij}^{2}-\sqrt{X_{ij}}\left(v_{ij}v_{ik}+v_{ik}v_{jk}-v_{ii}v_{jk}-v_{kk}v_{ij}-v_{ik}^{2}+v_{ii}v_{kk}\right)$
$\qquad-\sqrt{X_{ik}}\left(v_{ij}v_{ik}+v_{ij}v_{jk}-v_{jj}v_{ik}-v_{ii}v_{jk}-v_{ij}^{2}+v_{ii}v_{jj}\right)+\left(-v_{ij}-v_{ik}+v_{ii}+v_{jk}\right)\sqrt{X_{ij}X_{ik}}\Big{]}\,,$ (89)
$v_{(ij)(ki)}$ $=\frac{1}{(2v_{ij}-v_{ii}-v_{jj})(2v_{ik}-v_{ii}-v_{kk})}\Big{[}v_{ii}^{2}v_{jk}-v_{ii}v_{ij}v_{ik}-v_{ii}v_{ij}v_{jk}-v_{ii}v_{ik}v_{jk}+v_{ii}v_{jj}v_{kk}+v_{ij}v_{ik}^{2}$
$\qquad+v_{ij}^{2}v_{ik}-v_{jj}v_{ik}^{2}+v_{ij}v_{ik}v_{jk}-v_{kk}v_{ij}^{2}-\sqrt{X_{ij}}\left(v_{ij}v_{ik}+v_{ik}v_{jk}-v_{ii}v_{jk}-v_{kk}v_{ij}-v_{ik}^{2}+v_{ii}v_{kk}\right)$
$\qquad-\sqrt{X_{ik}}\left(-v_{ij}v_{ik}-v_{ij}v_{jk}+v_{jj}v_{ik}+v_{ii}v_{jk}+v_{ij}^{2}-v_{ii}v_{jj}\right)+\left(v_{ij}+v_{ik}-v_{ii}-v_{jk}\right)\sqrt{X_{ij}X_{ik}}\Big{]}\,,$ (90)
where $X_{ij}=v_{ij}^{2}-v_{ii}v_{jj}$
@@ Appendix B Full form of IBP relations
for $\text{for }\vec{\xi}_{1k}\text{: }0=$ $\left(-d+j_{1}+j_{2}+j_{3}+2\right)I_{j_{1}-1,j_{2},j_{3}}+\left(d-2j_{1}-j_{2}-j_{3}-2\right)I_{j_{1},j_{2},j_{3}}+j_{3}\,v_{13}\,I_{j_{1},j_{2},j_{3}+1}$
$+j_{2}\,v_{12}\,I_{j_{1},j_{2}+1,j_{3}}+j_{1}\,v_{11}\,I_{j_{1}+1,j_{2},j_{3}}-j_{3}\,I_{j_{1}-1,j_{2},j_{3}+1}-j_{2}\,I_{j_{1}-1,j_{2}+1,j_{3}}\,,$ (91)
for $\text{for }\vec{\xi}_{12}\text{: }0=$ $j_{3}\left(v_{23}-1\right)I_{j_{1}-1,j_{2},j_{3}+1}+j_{2}\left(v_{22}-1\right)I_{j_{1}-1,j_{2}+1,j_{3}}-j_{3}\left(v_{13}-1\right)I_{j_{1},j_{2}-1,j_{3}+1}$
@@
IBP relations for $\xi_{2k}$, $\xi_{3k}$, $\xi_{13}$, and $\xi_{23}$ follow by appropriate permutation of indices.
@@ Appendix C Recursion relations for three-denominator angular integrals
$I_{j_{1},j_{2},j_{3}}=\frac{1}{(1-j_{1})X_{123}}\Big{[}$ $-\left(v_{23}^{2}-v_{22}v_{33}\right)\left(-d+j+1\right)I_{j_{1}-2,j_{2},j_{3}}+\left(v_{13}v_{23}-v_{12}v_{33}\right)\left(-d+j+1\right)I_{j_{1}-1,j_{2}-1,j_{3}}$
$-\left(v_{13}v_{22}-v_{12}v_{23}\right)\left(-d+j+1\right)I_{j_{1}-1,j_{2},j_{3}-1}+[(v_{12}v_{23}-v_{13}v_{22})(d-j_{1}-j_{2}-2j_{3}-1)$
$+(v_{13}v_{23}-v_{12}v_{33})(d-j_{1}-2j_{2}-j_{3}-1)+\left(v_{22}v_{33}-v_{23}^{2}\right)(d-2j_{1}-j_{2}-j_{3})]I_{j_{1}-1,j_{2},j_{3}}$
@@
where $X_{123}$ is the Gram determinant defined in section IV.
There are also the symmetric identities with $1\leftrightarrow 2$ and $1\leftrightarrow 3$ applicable for $j_{2,3}\neq 1$. These identities allow for a systematic reduction of any $I_{j_{1},j_{2},j_{3}}$ with $j_{1},j_{2},j_{3}>0$ to the case where either $j_{1}=j_{2}=j_{3}=1$ or one of the indices becomes 00. From there, recursion relations for two denominators can be used for further reduction [5]. To deal with integrals with a negative index, we can use
@@
$\qquad+v_{23}\left(j_{2}\left(v_{12}-3\right)+\left(j_{3}-1\right)v_{13}-j_{1}\left(v_{12}+v_{13}+2\right)-v_{12}+2d-3j_{3}-6\right)$
$\qquad+v_{33}\left(\left(j_{1}-j_{2}+1\right)v_{12}-d+j_{1}+2j_{2}+j_{3}+3\right)$
$\qquad+v_{22}\left(d\left(v_{33}-1\right)+j_{1}v_{13}-j_{3}v_{13}-\left(2j_{1}+j_{2}+j_{3}+4\right)v_{33}+v_{13}+j_{1}+j_{2}+2j_{3}+3\right)]I_{j_{1}+1,j_{2},j_{3}}$
@@
which raises $j_{1}+j_{2}+j_{3}$ as well as $j_{1}$. Again there are the symmetric versions that raise $j_{2}$ respectively $j_{3}$. These can be used to increase any negative index to 00 again resulting in two-denominator integrals. We note that $Y_{23}$ has the geometrical interpretation $Y_{23}=4\mathbb{V}^{2}(\vec{v}_{2},\vec{v}_{3})$, where $\mathbb{V}(\vec{v}_{2},\vec{v}_{3})$ denotes the area of the triangle spanned by $\vec{v}_{2,3}$.

### v4

@@ 8.2 Single-massive integral
The quantities $Y_{123}^{(0,1)}$ are given in eqs. (40) respectively (41). As in the massless case, the result simplifies to a sum of two-denominator integrals in the case where the vectors $\vec{v}_{1,2,3}$ become linearly dependent. The pole part is identical to the result found in Smirnov:2024 , while the finite part agrees in the limit of small mass $v_{11}$ with the result calculated in Smirnov:2024 via expansion by regions.
8.3 Double- and triple-massive integral
The $\varepsilon$-expansions of the double- and triple-massive integral could be obtained by plugging the massless and single-massive results from eqs.(67) and (74) into eqs.(15) and (16). However, it is much more economical to first use eq. (25) to extract the part that is expressible in terms of known two-denominator integrals. Subsequently eqs.(15) and (16) need to be applied only to $I_{1,1,1}^{d=6}$. By absorbing a factor of $\sqrt{Y_{123}}$, we can apply the mass splitting in the form of (22) respectively (23). Therefore the order $\varepsilon$ term in the double- respectively triple-massive case is conveniently expressed in terms of the functions
$f_{Y}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})=$ $f_{Y}^{(0)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)}\right)-f_{Y}^{(1)}\left(v_{13},v_{1(12)},v_{3(12)},v_{11}\right)$
$-f_{Y}^{(1)}\left(v_{23},v_{2(21)},v_{3(21)},v_{22}\right),$ (79)
@@ 8.3 Double- and triple-massive integral
$f_{Y}^{(3)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22},v_{33})=f_{Y}^{(1)}\left(v_{3(12)},v_{3(21)},v_{(12)(21)},v_{33}\right)$
$-f_{Y}^{(0)}\left(v_{(12)(13)},v_{(12)(31)},v_{(13)(31)}\right)+f_{Y}^{(1)}\left(v_{1(12)},v_{1(13)},v_{(12)(13)},v_{11}\right)$
$+f_{Y}^{(1)}\left(v_{3(12)},v_{3(31)},v_{(12)(31)},v_{33}\right)-f_{Y}^{(0)}\left(v_{(21)(23)},v_{(21)(32)},v_{(23)(32)}\right)$
@@
$I_{1,1,1}^{(2)}$ $(v_{12},v_{13},v_{23},v_{11},v_{22};\varepsilon)=-\frac{\pi}{\varepsilon}\frac{1}{v_{13}v_{23}}+\frac{\pi}{2v_{12}v_{23}v_{13}-v_{11}v_{23}^{2}-v_{22}v_{13}^{2}}$
$\times\left[\Big(v_{13}-v_{12}-v_{23}+\frac{v_{11}v_{23}}{v_{13}}\Big)\log\frac{v_{11}}{v_{13}^{2}}\right.+\Big(v_{23}-v_{12}-v_{13}+\frac{v_{22}v_{13}}{v_{23}}\Big)\log\frac{v_{22}}{v_{23}^{2}}$
@@
$+\frac{\pi}{2v_{12}v_{23}v_{13}-v_{11}v_{23}^{2}-v_{22}v_{13}^{2}}$
$\times\left[\frac{1}{\sqrt{X_{12}}}\Big(v_{12}\left(v_{13}+v_{23}-v_{12}\right)-v_{11}v_{23}-v_{22}v_{13}+v_{11}v_{22}\Big)f_{1}^{(2)}(v_{12},v_{11},v_{22})\right$.
@@
$\left.\quad+2\sqrt{Y_{123}^{(2)}}\,f_{Y}^{(2)}(v_{12},v_{13},v_{23},v_{11},v_{22})\right]+\mathcal{O}\!\left(\varepsilon^{2}\right)\,.$ (84)
with $f_{Y}^{(2)}$ from eq. (79). Finally, using the kinematic quantities from section 4, the triple-massive three-denominator integral has the $\varepsilon$-expansion
$I_{1,1,1}^{(3)}$ $=\frac{\pi}{X_{123}^{(3)}}\left[\frac{X_{12\bar{3}}}{\sqrt{X_{12}}}\log\!\left(\frac{v_{12}+\sqrt{X_{12}}}{v_{12}-\sqrt{X_{12}}}\right)+\frac{X_{1\bar{2}3}}{\sqrt{X_{13}}}\log\!\left(\frac{v_{13}+\sqrt{X_{13}}}{v_{13}-\sqrt{X_{13}}}\right)\right$.
$\left.\qquad\qquad+\frac{X_{\bar{1}23}}{\sqrt{X_{23}}}\log\!\left(\frac{v_{23}+\sqrt{X_{23}}}{v_{23}-\sqrt{X_{23}}}\right)\right]$
$+\frac{\pi\,\varepsilon}{X_{123}^{(3)}}\left[\frac{X_{12\bar{3}}}{\sqrt{X_{12}}}f_{1}^{(2)}(v_{12},v_{11},v_{22})+\frac{X_{1\bar{2}3}}{\sqrt{X_{13}}}f_{1}^{(2)}(v_{13},v_{11},v_{33})+\frac{X_{\bar{1}23}}{\sqrt{X_{23}}}f_{1}^{(2)}(v_{23},v_{22},v_{33})\right$.
$\left.\qquad\qquad+2\sqrt{Y_{123}^{(3)}}\,f_{Y}^{(3)}\right]+\mathcal{O}\!\left(\varepsilon^{2}\right)$ (85)
with $f_{Y}^{(3)}$ from eq. (80). For the double- and triple-massive integral, poles and small-mass limit of the finite part are also in agreement with the results from Smirnov:2024 .
9 Beyond three denominators
If we consider a situation in $d=4-2\varepsilon$ dimensions where all vectors $v_{i}$ are confined to the physical $4$-dimensional subspace – an assumption applicable if the $v_{i}$ correspond to momenta of observed particles – we can reduce angular integrals with an arbitrary number of denominators raised to integer powers to the three-denominator case presented in this paper. This is possible since in $4$ dimensions, i.e. three spatial dimensions, no more than three $\vec{v}_{i}$ can be linearly dependent. Hence for any $n>3$ and vectors $v_{1},\dots,v_{n}$ there are constants $\lambda_{i}$ such that $\lambda\equiv\sum_{i=1}^{n}\lambda_{i}\neq 0$ but $0=\sum_{i=1}^{n}\lambda_{i}\vec{v}_{i}$. By partial fractioning it holds that Lyubovitskij:2021
@@ 10 Conclusion
In parallel work, Taushif Ahmed, Syed Mehedi Hasan, and Andreas Rapakoulias independently calculated the angular integral with three denominators using Mellin-Barnes integrals Ahmed:2024 . Preliminary numerical comparison at selected phase space points showed agreement.
Acknowledgements.
@@ Appendix A Scalar products for mass reduction formula
The scalar products involving the massless auxiliary vectors appearing in section 3 are given by ($i,j,k\in\{1,2,3\}$)
@@
$v_{(ij)(ik)}$ $=\frac{1}{(2v_{ij}-v_{ii}-v_{jj})(2v_{ik}-v_{ii}-v_{kk})}$
$\times$ $\Big[v_{ii}^{2}v_{jk}-v_{ii}v_{ij}v_{ik}-v_{ii}v_{ij}v_{jk}-v_{ii}v_{ik}v_{jk}+v_{ii}v_{jj}v_{kk}+v_{ij}v_{ik}^{2}+v_{ij}^{2}v_{ik}-v_{jj}v_{ik}^{2}$
$+v_{ij}v_{ik}v_{jk}-v_{kk}v_{ij}^{2}-\sqrt{X_{ij}}\left(v_{ij}v_{ik}+v_{ik}v_{jk}-v_{ii}v_{jk}-v_{kk}v_{ij}-v_{ik}^{2}+v_{ii}v_{kk}\right)$
$-\sqrt{X_{ik}}\left(v_{ij}v_{ik}+v_{ij}v_{jk}-v_{jj}v_{ik}-v_{ii}v_{jk}-v_{ij}^{2}+v_{ii}v_{jj}\right)$
$+\left(-v_{ij}-v_{ik}+v_{ii}+v_{jk}\right)\sqrt{X_{ij}X_{ik}}\Big]\,$, (90)
$v_{(ij)(ki)}$ $=\frac{1}{(2v_{ij}-v_{ii}-v_{jj})(2v_{ik}-v_{ii}-v_{kk})}$
$\times$ $\Big[v_{ii}^{2}v_{jk}-v_{ii}v_{ij}v_{ik}-v_{ii}v_{ij}v_{jk}-v_{ii}v_{ik}v_{jk}+v_{ii}v_{jj}v_{kk}+v_{ij}v_{ik}^{2}+v_{ij}^{2}v_{ik}-v_{jj}v_{ik}^{2}$
$\qquad+v_{ij}v_{ik}v_{jk}-v_{kk}v_{ij}^{2}-\sqrt{X_{ij}}\left(v_{ij}v_{ik}+v_{ik}v_{jk}-v_{ii}v_{jk}-v_{kk}v_{ij}-v_{ik}^{2}+v_{ii}v_{kk}\right)$
$\qquad-\sqrt{X_{ik}}\left(-v_{ij}v_{ik}-v_{ij}v_{jk}+v_{jj}v_{ik}+v_{ii}v_{jk}+v_{ij}^{2}-v_{ii}v_{jj}\right)$
$+\left(v_{ij}+v_{ik}-v_{ii}-v_{jk}\right)\sqrt{X_{ij}X_{ik}}\Big]\,$, (91)
where $X_{ij}=v_{ij}^{2}-v_{ii}v_{jj}$ .
@@ Appendix B Full form of IBP relations
$\text{for }\vec{\xi}_{1k}\text{: }0=\left(-d+j_{1}+j_{2}+j_{3}+2\right)I_{j_{1}-1,j_{2},j_{3}}+\left(d-2j_{1}-j_{2}-j_{3}-2\right)I_{j_{1},j_{2},j_{3}}$
$\;\;+j_{3}\,v_{13}\,I_{j_{1},j_{2},j_{3}+1}+j_{2}\,v_{12}\,I_{j_{1},j_{2}+1,j_{3}}+j_{1}\,v_{11}\,I_{j_{1}+1,j_{2},j_{3}}-j_{3}\,I_{j_{1}-1,j_{2},j_{3}+1}-j_{2}\,I_{j_{1}-1,j_{2}+1,j_{3}}\,$, (92)
$\text{for }\vec{\xi}_{12}\text{: }0=j_{3}\left(v_{23}-1\right)I_{j_{1}-1,j_{2},j_{3}+1}+j_{2}\left(v_{22}-1\right)I_{j_{1}-1,j_{2}+1,j_{3}}-j_{3}\left(v_{13}-1\right)I_{j_{1},j_{2}-1,j_{3}+1}$
@@
IBP relations for $\vec{\xi}_{2k}$, $\vec{\xi}_{3k}$, $\vec{\xi}_{13}$, and $\vec{\xi}_{23}$ follow by appropriate permutation of indices.
@@ Appendix C Recursion relations for three-denominator angular integrals
$I_{j_{1},j_{2},j_{3}}=\frac{1}{(1-j_{1})X_{123}}\Big[-\left(v_{23}^{2}-v_{22}v_{33}\right)\left(-d+j+1\right)I_{j_{1}-2,j_{2},j_{3}}$
$+\left(v_{13}v_{23}-v_{12}v_{33}\right)\left(-d+j+1\right)I_{j_{1}-1,j_{2}-1,j_{3}}-\left(v_{13}v_{22}-v_{12}v_{23}\right)\left(-d+j+1\right)I_{j_{1}-1,j_{2},j_{3}-1}$
$+[(v_{12}v_{23}-v_{13}v_{22})(d-j_{1}-j_{2}-2j_{3}-1)+(v_{13}v_{23}-v_{12}v_{33})(d-j_{1}-2j_{2}-j_{3}-1)$
$\quad+\left(v_{22}v_{33}-v_{23}^{2}\right)(d-2j_{1}-j_{2}-j_{3})]I_{j_{1}-1,j_{2},j_{3}}$
@@
where $X_{123}$ is the Gram determinant defined in section 4.
There are also the symmetric identities with $1\leftrightarrow 2$ and $1\leftrightarrow 3$ applicable for $j_{2,3}\neq 1$. These identities allow for a systematic reduction of any $I_{j_{1},j_{2},j_{3}}$ with $j_{1},j_{2},j_{3}>0$ to the case where either $j_{1}=j_{2}=j_{3}=1$ or one of the indices becomes 0. From there, recursion relations for two denominators can be used for further reduction Lyubovitskij:2021 . To deal with integrals with a negative index, we can use
@@
$\quad+v_{23}\left(j_{2}\left(v_{12}-3\right)+\left(j_{3}-1\right)v_{13}-j_{1}\left(v_{12}+v_{13}+2\right)-v_{12}+2d-3j_{3}-6\right)$
$\quad+v_{33}\left(\left(j_{1}-j_{2}+1\right)v_{12}-d+j_{1}+2j_{2}+j_{3}+3\right)+v_{22}\left(d\left(v_{33}-1\right)+j_{1}v_{13}-j_{3}v_{13}\right$.
$\quad\left.\quad-\left(2j_{1}+j_{2}+j_{3}+4\right)v_{33}+v_{13}+j_{1}+j_{2}+2j_{3}+3\right)]I_{j_{1}+1,j_{2},j_{3}}$
@@
which raises $j_{1}+j_{2}+j_{3}$ as well as $j_{1}$. Again there are the symmetric versions that raise $j_{2}$ respectively $j_{3}$. These can be used to increase any negative index to 0 again resulting in two-denominator integrals. We note that $Y_{23}$ has the geometrical interpretation $Y_{23}=4\mathbb{V}^{2}(\vec{v}_{2},\vec{v}_{3})$, where $\mathbb{V}(\vec{v}_{2},\vec{v}_{3})$ denotes the area of the triangle spanned by $\vec{v}_{2,3}$.